OpenStudy (anonymous):

Please show me the work. 1. (4x/9) + (x-10/9) 2. (t^2+6t/t-1) -(7/t-1) 3. (3/x)+ (13/x-11) 4. (5/4x^2) - (y/14xz) 5. (x^2+2x/12x+54) - (3-x/8x+36) 6. (2/x^28x+15) +(1/x^2+11x+30)

OpenStudy (anonymous):

1. (4x/9) + (x-10/9) =(4/9)x +x - (10/9) =(1+ 4/9)x - 10/9 =13x/9 -10/9

OpenStudy (anonymous):

Thank You.. Will you please do the same for #2,3,4,5,6 :)